1gr of graphite is burnt in a bomb calorimeter in excess of oxygen at 298k and 1atm pressure according to the equation
C(graphite) +O2(g) = CO2(g)
During the reaction temperature rises from298Kto299K.If the heat capacity of the bomb calormetercais 20.7KJ/mole . What is the enthalpy change for the above reaction at 298K and 1atm.

Grade:11

1 Answers

Arun
5 years ago

Rise in temperature,  ΔT  = 299-298 = 1 k

Heat capacity of calorimeter ,  Cv  = 20.7 kJ/mol (given)

so, Heat absorbed by calorimeter  = C× ΔT = 20.1 × 1= 20.1kJ

This is the heat evolved in combustion of 1g graphite

So, heat evolved in the combustion of 1 mole of graphite

i.e. 12 g of graphite  = 20.7 ×  12  = 284.4kJ

Since heat is evolved and vessel is closed

so, enthalpy change of the reaction  = - 284.4kJ/ mol